1. A local grocery store purchases a dozen of eggs for P60. 00 and sell it for P84. 00. Compute for the:


a. Mark-up

b. Mark-down

c. Gross margin

d. Gross margin ratio


3. After a reduction of trade discount rate of 25% of the merchandise purchased, the invoice price was P13, 000.00. what was the list price?


4. A television is listed at P333,000.00 less discount of 5%, 7% and 9%. Fins the net price.


5. Two suppliers offer identical sets of kitchen wares at the same list price. However, company ABC gives trade discount series of 20%, 15% and 5%, while company DEF offers 25%, 10% and 54%. Which is a better discount for the buyer?


1
Expert's answer
2022-01-10T09:56:50-0500

1.

a. Mark-up is (84 - 60)/60 = 0.4 or 40%.

b. There is no mark-down.

c. Gross margin is 84 - 60 = P24.

d. Gross margin ratio is 24/60 = 0.4 or 40%.


3. The list price is:

13,000/(1 - 0.25) = 17,333.33.


4. The net price is:

333,000×(1 - 0.05 - 0.07 - 0.09) = 263,070.


5. DEF provides a better discount for the buyer.